OpenStudy (anonymous):

please help nee this today there r several forms for the equation of a line. each form communicates different information. for instance, from the point-slope form,you can determine a point and the slope of a line. write an equation in slope-intercept, point-slope, or standard form for the line with the given information. explain why u chose the form u used. a. passes through (-1,4)and(-5,2) b. slope 2, y-intercept -4 c.has an x-intercept of 6 and a y-intercept of 3 d. passes through (1,2) with slope -5/3

OpenStudy (anonymous):

b. y = 2x -4......coz slope intercept coz slope nd intercept are given

OpenStudy (anonymous):

c. x/6 + y/ 3 =1...intercept form coz intercepts r given

OpenStudy (anonymous):

d. use slope intercept and a use two point form

OpenStudy (anonymous):

a. passes through points, u use the gradient formula and find an equation formula for this \[M=\frac{2-4}{-5+2} \] M = -2/ -3 M = 2/3 sub that gradient and a point given (-1,4) into the find an equation formula y - 4 = 2/3 (x - -1) y - 4 = 2/3 (x+1) y - 4 = 2/3x + 2/3 y = 2/3x + 2/3 + 4 y = 2/3x + 14/3

OpenStudy (anonymous):

aww thanks:)

OpenStudy (anonymous):

for c. you can use two point form just like in a but there is also this equation of a line: \[\large \frac{x}{a}+\frac{y}{b}=1 \] where a and b are you x and y intercepts, respectively...

OpenStudy (anonymous):

thanks